OpenStudy (anonymous):

is parallel to y=-4x-9 and passes through the point (0,2). write in slope intercept form

OpenStudy (anonymous):

Find the slope(m) from the given equation - \[ y=-4x-9 \] Then use this equation , \[(y-y1)=m(x-x1) \] (0,2)=(x1,y1)

OpenStudy (anonymous):

You know the slope is 4 since they are parallel. So take the point, and the slope, and fill it in for y=mx+b So 2=4(0)+b and solve for b. which is 2. so y=4x+2.. Hopefully this helps...